# Bayesian Component Separation for DESI LAE Automated Spectroscopic Redshifts and Photometric Targeting

### v1

@@ Abstract
Lyman Alpha Emitters (LAEs) are valuable high-redshift cosmological probes traditionally identified using specialized narrow-band photometric surveys. In ground-based spectroscopy, it can be difficult to distinguish the sharp LAE peak from residual sky emission lines using automated methods, leading to misclassified redshifts. We present a Bayesian spectral component separation technique to automatically determine spectroscopic redshifts for LAEs while marginalizing over sky residuals. We use visually inspected spectra of LAEs obtained using the Dark Energy Spectroscopic Instrument (DESI) to create a data-driven prior and can determine redshift by jointly inferring sky residual, LAE, and residual components for each individual spectrum. We demonstrate this method on 910 spectroscopically observed DESI LAE candidate spectra and determine their redshifts with 90% accuracy when validated against visually inspected redshifts. Using the value from our pipeline as a proxy for detection confidence, we then explore potential survey design choices and implications for targeting LAEs with medium-band photometry. This method allows for scalability and accuracy in determining redshifts from DESI spectra, and the results provide recommendations for LAE targeting in anticipation of future high-redshift spectroscopic surveys.
1 Introduction
@@
Large-scale observational studies of LAEs have been key to realizing their potential as cosmological probes. LAEs have been targeted mainly through spectroscopy and narrow-band photometry, and have been observed from the ground and space, including with Keck (Spinrad et al., 1997; Dey et al., 1998; Hu et al., 1998, 1999; Stern et al., 2000; Manning et al., 2000; Cooper et al., 2024), Very Large Telescope (Kudritzki et al., 2000), Hubble Space Telescope (Pascarelle et al., 1998), Dark Energy Camera (DECam; Lee et al., 2024), Subaru Hyper-Suprime Cam (Kawanomoto et al., 2018; Aihara et al., 2019; Kikuta et al., 2023), James Webb Space Telescope (Kumari et al., 2024), and most recently, the Dark Energy Spectroscopic Instrument (DESI; Dey et al., 2025).
The Dark Energy Spectroscopic Instrument (DESI) has collected spectra of millions of galaxies for precise cosmological measurements (DESI Collaboration et al., 2016a, 2022, 2025). While LAEs are not one of the primary targets, they have been observed by DESI through secondary and tertiary targeting programs, providing spectroscopic follow-up for LAE candidates identified from photometric surveys (Dey et al., 2025; Raichoor et al., 2025). After DESI completes its initial five-year run, the planned follow-up program (DESI-2) will focus on higher-redshift targets and aims to observe 1.5 million LAEs (Schlegel et al., 2022; Ravi et al., 2024).
The current DESI spectroscopic data analysis pipeline (Bailey et al., 2025) does not automatically identify or determine redshifts for LAEs, which are instead currently done by visual inspection. In anticipation of LAEs becoming a primary target for DESI-2, it is important to have reliable, automated software pipelines to identify and characterize them. This requires the use of principled data methods that maximize interpretability while being computationally efficient.
@@
An outline of this paper is as follows: in Section 2 we describe the spectra being used and how the objects were targeted. In Section 3 we describe in detail the methods we use, how we generate our priors for each component, and how they are applied to the data. The results of the redshift pipeline on real and simulated data and applications to LAE targeting are presented in Section 4 and discussed further in Section 5 before we conclude in Section 6.
2 Data
2.1 DESI
DESI is an optical multi-object fiber spectrograph located on the Nicholas U. Mayall 4-meter Telescope at the Kitt Peak National Observatory in Arizona (Levi et al., 2013; DESI Collaboration et al., 2016b, 2022). It uses robotic fiber positioners that allow it to collect 5,000 high-quality spectra simultaneously (DESI Collaboration et al., 2016c; Silber et al., 2023; Miller et al., 2024). DESI spectra span Å with between 2000 and 5000, and are binned linearly to Å pixels (DESI Collaboration et al., 2016a; Poppett et al., 2024). For more information about survey validation and the Early Data Release, see DESI Collaboration et al. (2024a, b); for Data Release 1, see DESI Collaboration et al. (2025), and for key cosmological results, see DESI Collaboration et al. (2024c, d, e, f, g, h).
The four main galaxy samples targeted by DESI are bright galaxies, luminous red galaxies (LRGs), emission line galaxies (ELGs), and QSOs (Myers et al., 2023; Schlafly et al., 2023). The DESI Collaboration has an extensive spectroscopic data analysis pipeline that includes automated redshift determination for each of these four main targeted object types (Guy et al., 2023; Bailey et al., 2025; Brodzeller et al., 2023). The current automated redshift determination pipeline, Redrock, uses a Principal Component Analysis (PCA)-based approach and has templates for stars, galaxies, and quasars (Guy et al., 2023; Bailey et al., 2025). DESI has also implemented a significant visual inspection (VI) campaign for tens of thousands of galaxy spectra, where collaboration members look at individual spectra and report a redshift estimate, as well as provide written comments about its spectral type (Lan et al., 2023). Currently, the Redrock redshift pipeline does not yield accurate redshifts for the LAE samples, and instead redshift measurement has been done via visual inspection.
@@ 2.2 The LAE Samples
For this study, we used two different samples of LAEs targeted by DESI: one selected from the One-hundred deg2 DECam Imaging in Narrowbands (ODIN) survey and the second selected from archival Subaru Suprime-Cam intermediate-band imaging data. We treat the narrow-band targeted ODIN LAE candidates, which are a larger sample size, as known LAEs, and use them to create our data-driven LAE prior. This prior is then applied to infer the presence of an LAE and its redshift in spectra of the Subaru LAE candidates, which we use as our test set. The details of the target selection and spectroscopy of these LAE samples are described in more detail elsewhere (see Dey et al., 2025; Raichoor et al., 2025), but we provide a brief description of these data here.
2.2.1 ODIN LAEs
@@
The One-hundred deg2 DECam Imaging in Narrowbands (ODIN) survey is a photometric survey specifically designed to find potential LAEs by identifying targets with narrow-band excess flux (Lee et al., 2024). It uses the Dark Energy Camera (DECam) at the prime focus of the 4m Victor Blanco Telescope located at the Cerro Tololo Inter-American Observatory (CTIO) in Chile. ODIN uses three 70–100 Å FWHM narrow-band filters (N419, N501, and N673, seen in Figure 1) to image seven fields covering a total area of to depths of 25.5 – 25.9 mag (Firestone et al., 2024). Dey et al. (2025) used the ODIN imaging data in the COSMOS and XMM fields in conjunction with archival imaging from the Subaru Hyper-SuprimeCam Subaru Strategic Program survey (Aihara et al., 2018, 2019, 2022) and imaging from the DESI Legacy Imaging Surveys (Dey et al., 2019) in regions not covered by the HSC imaging. Photometric catalogs of these data were created using Tractor (Lang et al., 2016) and LAE candidates were selected in the COSMOS and XMM fields (see Dey et al. (2025) for details). DESI observations of 11,570 of these candidates were carried out during 2022 and 2023. All the spectra were processed by the DESI spectroscopic reduction pipeline, and all spectra were visually inspected by members of the DESI and ODIN collaboration to measure redshifts. The visual inspection results in a redshift estimate (typically determined from the peak of the Ly emission line) and a quality flag (‘VI_QUALITY’) which ranges from 0 to 4, with higher numbers characteristic of spectra with more well-detected or multiple lines and, thus, more confident redshifts. The visual inspections also provide labels for the spectra, identifying LAEs, QSOs, and low-redshift galaxy interlopers (see Dey et al. (2025) for further details). From these observations, we selected a subset of 5,436 non-QSO LAEs at redshifts with robust visually-confirmed redshifts (i.e., VI_QUALITY ); this subset, which contains objects selected with all three ODIN narrow bands, is used for constructing the data-driven priors.
2.2.2 Subaru Suprime-Cam LAEs
@@
As in the case with ODIN, photometric catalogs of these data were created in combination with broad-band imaging data from the Hyper-SuprimeCam SSP (which cover the 2 deg2 footprint) and selected LAE candidates as described in Raichoor et al. (2025). DESI targeted 1860 of these candidates during 2023, and the pipeline-processed spectra of all targeted sources were visually inspected in the same manner as described for the ODIN sources. Of these, a subset of 910 LAEs with robust visually inspected redshifts between 2 and 4 were selected to be the test sample for this paper.
3 Method Overview
@@
The technique used in this work, Marginalized Analytic Dataspace Gaussian Inference for Component Separation (MADGICS), is a covariance-based Bayesian component separation technique (Saydjari et al., 2023, 2025a, 2025b). MADGICS has previously been used to pinpoint the wavelength and width of diffuse interstellar bands in Gaia RVS spectra (Saydjari et al., 2023) and also to improve the precision of stellar radial velocity measurements in APOGEE (Saydjari et al., 2025a).
@@
This should be negative, as including an LAE component can only increase the goodness of fit.
3.1 Prior generation
3.1.1 Sky residual prior
@@
We then use these rescaled spectra to create a new data-driven covariance matrix. From there, the top 50 eigenvectors, which were determined to preserve most of the variance of the sky residual spectra, are used to create a low-rank approximation of the covariance matrix. Thus, we express the sky residual covariance , where is the matrix of the top 50 eigenvectors scaled by the square root of their associated eigenvalues. This serves as our prior for the sky residual component.
3.1.2 LAE prior
To create the LAE prior, we use DESI spectra from ODIN LAE candidates with VI redshifts between 2 – 4.5. Here we consider the VI redshifts to be ground truth, although in practice there is likely a small error rate for both LAE identification and redshift precision. In this work, we only use LAE target spectra with an average “VI_QUALITY” score 2.5 (on a scale from 0 to 4), and which did not have any mention of the phrase “QSO” in the VI comments.
These spectra interpolated to the same log-spacing as the sky residual spectra (). They are then scaled as in Equation 8 and decomposed into sky residual and residual components using Equation 2. We use the residual component from this decomposition so as to not include sky residual lines in the prior covariance matrix. To ensure that we use high SNR spectra, we calculate the variance of the flux values of the spectrum and include only those with variances below the 75th percentile of the whole population to exclude overly noisy spectra, leaving 4,125 spectra. Given the visually inspected redshifts, we shift the remaining spectra to all be at , the central redshift of the N419 ODIN filter.
@@ 3.1.2 LAE prior
For these spectra, redshifts are determined based on the observed Ly emission line, which can be offset from the systemic redshift of the galaxy by 100 – 300 km/s; this should not significantly impact cosmological analyses done at large length scales, but may impact small-scale analyses.
3.2 Spectroscopic redshift determination
@@
At each redshift step, we calculate the as in Equation 6 and find the redshift that minimizes this profile, which indicates that including an LAE component at that redshift improves the fit the most. The also indicates a level of confidence or SNR in our detection; higher indicates a more robust detection of an LAE. To achieve a more precise redshift estimate, we do an additional, finer scan along 5 pixels around this best redshift value in steps of 0.1 pixels, and assign the target the minimum redshift from this final scan. Averaged over the 910 Subaru targets, it took seconds to determine the redshift for one target spectrum.
@@ 4 Results & Validation
4.1 Redshift determination for Subaru LAE targets
We use the priors described in Section 3 to decompose and determine redshifts for DESI spectra of 910 LAE candidates targeted with the Subaru that have VI redshifts between 2 and 4. We evaluate our method on a held-out test set, meaning that none of the spectra for which we show results were used to create the LAE prior covariance matrix.
Figure 4 shows the (calibrated as described in Section 4.2.3) as a function of redshift residual (new - VI ) for this test set. There is generally good agreement between the redshifts determined by our pipeline and those from visual inspection: 836 (91.9%) of the 910 evaluated spectra were determined by our method to have a redshift within 0.005 of the VI. Most disagreements between the two methods occur at low , which would indicate low SNR or an uncertain detection. Several factors could contribute to our pipeline determining redshifts outside of this threshold: some of the sources are Lyman Break Galaxies (LBGs) or other objects which do not have a strong Ly emission line; other objects show strongly double-peaked Ly , or very weak lines; and some are erroneous redshifts reported by the visual inspection.
To create a binary classifier (LAE or not LAE), one could impose a threshold value, above which any target is classified as being an LAE. Figure 5 shows the purity and completeness of the LAE sample as a function of the potential calibrated threshold value, where a “correct” detection is a target that has a pipeline redshift within 0.005 of its VI redshift. Note that we are evaluating our method on a pre-selected sample of solely LAE targets with no known non-LAEs, so a high purity across all thresholds is to be expected. Figures 4 and 5 suggest that a value of 25 could be an effective threshold value for classifying an object as an LAE, which roughly corresponds to a 5 cut.
4.2 Injection-recovery tests
4.2.1 Uniform noise

### v2

@@ Abstract
Lyman Alpha Emitters (LAEs) are valuable high-redshift cosmological probes traditionally identified using specialized narrow-band photometric surveys. In ground-based spectroscopy, it can be difficult to distinguish the sharp LAE peak from residual sky emission lines using automated methods, leading to misclassified redshifts. We present a Bayesian spectral component separation technique to automatically determine spectroscopic redshifts for LAEs while marginalizing over sky residuals. We use visually inspected spectra of LAEs obtained using the Dark Energy Spectroscopic Instrument (DESI) to create a data-driven prior and can determine redshift by jointly inferring sky residual, LAE, and residual components for each individual spectrum. We demonstrate this method on 881 spectroscopically observed DESI LAE candidate spectra and determine their redshifts with 90% accuracy when validated against visually inspected redshifts. Using the value from our pipeline as a proxy for detection confidence, we then explore potential survey design choices and implications for targeting LAEs with medium-band photometry. This method allows for scalability and accuracy in determining redshifts from DESI spectra, and the results provide recommendations for LAE targeting in anticipation of future high-redshift spectroscopic surveys.
I Introduction
@@
Large-scale observational studies of LAEs have been key to realizing their potential as cosmological probes. LAEs have been targeted mainly through spectroscopy and narrow-band photometry, and have been observed from the ground and space, including with Keck (Spinrad et al., 1997; Dey et al., 1998; Hu et al., 1998, 1999; Stern et al., 2000; Manning et al., 2000; Cooper et al., 2024), Very Large Telescope (Kudritzki et al., 2000), Hubble Space Telescope (Pascarelle et al., 1998), Dark Energy Camera (DECam; Lee et al., 2024), Subaru Hyper-Suprime Cam (Kawanomoto et al., 2018; Aihara et al., 2019; Kikuta et al., 2023), James Webb Space Telescope (Kumari et al., 2024), and most recently, the Dark Energy Spectroscopic Instrument (DESI; Dey et al., 2026).
The Dark Energy Spectroscopic Instrument (DESI) has collected spectra of millions of galaxies for precise cosmological measurements (DESI Collaboration et al., 2016a, 2022, 2025). While LAEs are not one of the primary targets, they have been observed by DESI through secondary and tertiary targeting programs, providing spectroscopic follow-up for LAE candidates identified from photometric surveys (Dey et al., 2026; Raichoor et al., 2026). After DESI completes its initial five-year run, the planned follow-up program (DESI-2) will focus on higher-redshift targets and aims to observe 1.5 million LAEs (Schlegel et al., 2022; Ravi et al., 2024).
The current DESI spectroscopic data analysis pipeline (Bailey et al., 2026) does not automatically identify or determine redshifts for LAEs, which are instead currently done by visual inspection. In anticipation of LAEs becoming a primary target for DESI-2, it is important to have reliable, automated software pipelines to identify and characterize them. This requires the use of principled data methods that maximize interpretability while being computationally efficient.
@@
An outline of this paper is as follows: in Section II we describe the spectra being used and how the objects were targeted. In Section III we describe in detail the methods we use, how we generate our priors for each component, and how they are applied to the data. The results of the redshift pipeline on real and simulated data and applications to LAE targeting are presented in Section IV and discussed further in Section V before we conclude in Section VI.
II Data
II.1 DESI
DESI is an optical multi-object fiber spectrograph located on the Nicholas U. Mayall 4-meter Telescope at the Kitt Peak National Observatory in Arizona (Levi et al., 2013; DESI Collaboration et al., 2016b, 2022). It uses robotic fiber positioners that allow it to collect 5,000 high-quality spectra simultaneously (DESI Collaboration et al., 2016c; Silber et al., 2023; Miller et al., 2024). DESI spectra span Å with between 2000 and 5000, and are binned linearly to Å pixels (DESI Collaboration et al., 2016a; Poppett et al., 2024). For more information about survey validation and the Early Data Release, see DESI Collaboration et al. (2024a, b); for Data Release 1, see DESI Collaboration et al. (2025), and for key cosmological results, see the DESI key papers (DESI Collaboration et al., 2024c, d, e, f, g, h).
The four main galaxy samples targeted by DESI are bright galaxies, luminous red galaxies (LRGs), emission line galaxies (ELGs), and QSOs (Myers et al., 2023; Schlafly et al., 2023). The DESI Collaboration has an extensive spectroscopic data analysis pipeline that includes automated redshift determination for each of these four main targeted object types (Guy et al., 2023; Bailey et al., 2026; Brodzeller et al., 2023). The current automated redshift determination pipeline, Redrock, uses a Principal Component Analysis (PCA)-based approach and has templates for stars, galaxies, and quasars (Guy et al., 2023; Bailey et al., 2026). DESI has also implemented a significant visual inspection (VI) campaign for tens of thousands of galaxy spectra, where collaboration members look at individual spectra and report a redshift estimate, as well as provide written comments about its spectral type (Lan et al., 2023). Currently, the Redrock redshift pipeline does not yield accurate redshifts for the LAE samples, and instead redshift measurement has been done via visual inspection.
@@ II.2 The LAE Samples
For this study, we used two different samples of LAEs targeted by DESI: one selected from the One-hundred deg2 DECam Imaging in Narrowbands (ODIN) survey and the second selected from archival Subaru Suprime-Cam intermediate-band imaging data. We treat the narrow-band targeted ODIN LAE candidates, which are a larger sample size, as known LAEs, and use them to create our data-driven LAE prior. This prior is then applied to infer the presence of an LAE and its redshift in spectra of the Subaru LAE candidates, which we use as our test set. The details of the target selection and spectroscopy of these LAE samples are described in more detail elsewhere (see Dey et al., 2026; Raichoor et al., 2026), but we provide a brief description of these data here.
II.2.1 ODIN LAEs
@@
The One-hundred deg2 DECam Imaging in Narrowbands (ODIN) survey is a photometric survey specifically designed to find potential LAEs by identifying targets with narrow-band excess flux (Lee et al., 2024). It uses the Dark Energy Camera (DECam) at the prime focus of the 4m Victor Blanco Telescope located at the Cerro Tololo Inter-American Observatory (CTIO) in Chile. ODIN uses three 70–100 Å FWHM narrow-band filters (N419, N501, and N673, seen in Figure 1) to image seven fields covering a total area of to depths of 25.5 – 25.9 mag (Firestone et al., 2024). Dey et al. (2026) used the ODIN imaging data in the COSMOS and XMM fields in conjunction with archival imaging from the Subaru Hyper-SuprimeCam Subaru Strategic Program survey (Aihara et al., 2018, 2019, 2022) and imaging from the DESI Legacy Imaging Surveys (Dey et al., 2019) in regions not covered by the HSC imaging. New photometric catalogs of these data were created using Tractor (Lang et al., 2016) and LAE candidates were selected in the COSMOS and XMM fields (see Dey et al. (2026) for details). DESI observations of 11,634 of these candidates were carried out during 2022 and 2023.
All the spectra were processed by the DESI spectroscopic reduction pipeline, and all spectra were visually inspected by members of the DESI and ODIN collaboration to measure redshifts. The visual inspection results in a redshift estimate (typically determined from the peak of the Ly emission line) and a quality flag (‘VI_QUALITY’) which ranges from 0 to 4. Very briefly, the rubric used classifies quality as: “0” for no detected flux; “1” for some weak flux, perhaps continuum, but no redshift estimate; “2” for a weak feature detected at low signal-to-noise and a tentative redshift; “3” for a single line redshift (which is reliable when the line is Ly) by a break, and with no other significant lines in the DESI spectrum; or “4” for a firm redshift where two or more lines are detected. Each spectrum was inspected by at least two individuals, with the quality flag averaged between them.
The visual inspections also provide labels for the spectra, identifying LAEs, QSOs, and low-redshift galaxy interlopers (see Dey et al. (2026) for further details). QSOs were identified by visual inspectors likely based on the presence of broad emission lines and were noted in the VI comments; 763 objects visually identified as QSOs are excluded from this sample. From these observations and associated labels, we selected a subset of 5,341 non-QSO LAEs at redshifts with robust visually-confirmed redshifts (i.e., VI_QUALITY ). This subset, which contains objects selected with all three ODIN narrow bands, is used for constructing the data-driven priors.
II.2.2 Subaru Suprime-Cam LAEs
@@
As in the case with ODIN, photometric catalogs of these data were created in combination with broad-band imaging data from the Hyper-SuprimeCam SSP (which cover the 2 deg2 footprint) and selected LAE candidates as described in Raichoor et al. (2026). DESI targeted 1860 of these candidates during 2023, and the pipeline-processed spectra of all targeted sources were visually inspected in the same manner as described for the ODIN sources. Of these, a subset of 881 LAEs without visually identified QSO contamination with robust visually inspected redshifts between 2 and 4 were selected to be the test sample for this paper.
III Method Overview
@@
The technique used in this work, Marginalized Analytic Dataspace Gaussian Inference for Component Separation (MADGICS), is a covariance-based Bayesian component separation technique (Saydjari et al., 2023, 2025, 2026). MADGICS has previously been used to pinpoint the wavelength and width of diffuse interstellar bands in Gaia RVS spectra (Saydjari et al., 2023) and also to improve the precision of stellar radial velocity measurements in APOGEE (Saydjari et al., 2025).
@@
This should be negative, as including an LAE component can only increase the goodness of fit, yielding a lower value.
III.1 Prior generation
III.1.1 Sky residual prior
@@
We then use these rescaled spectra to create a new data-driven covariance matrix. From there, the top 50 eigenvectors222We also did the analysis with 25 and 100 eigenvectors, and found that performance did not improve beyond 50. We are incentivized to minimize the number of eigenvectors to optimize computational speed. are used to create a low-rank approximation of the covariance matrix. Thus, we express the sky residual covariance , where is the matrix of the top 50 eigenvectors scaled by the square root of their associated eigenvalues. This serves as our prior for the sky residual component.
III.1.2 LAE prior
To create the LAE prior, we use DESI spectra from ODIN LAE candidates with VI redshifts between 2.3 – 4.5. Here we consider the VI redshifts to be ground truth, although in practice there is likely a small error rate for both LAE identification and redshift precision. In this work, we only use LAE target spectra with an average “VI_QUALITY” score 2.5 (on a scale from 0 to 4), and which were not visually identified as QSOs, as discussed in Section II.2.1.
These spectra interpolated to the same log-spacing as the sky residual spectra (). They are then scaled as in Equation 8 and decomposed into sky residual and residual components using Equation 2. We use the residual component from this decomposition so as to not include sky residual lines in the prior covariance matrix. To ensure that we use high SNR spectra, we calculate the variance of the flux values of the spectrum and include only those with variances below the 75th percentile of the whole population to exclude overly noisy spectra, leaving 4,005 spectra. Given the visually inspected redshifts, we shift the remaining spectra to all be at , the central redshift of the N419 ODIN filter.
@@ III.1.2 LAE prior
For these spectra, redshifts are determined based on the observed Ly emission line, which can be offset from the systemic redshift of the galaxy by 100 – 300 km/s (Verhamme et al., 2018; Steidel et al., 2018; Pahl et al., 2021); this should not significantly impact cosmological analyses done at large length scales, but may impact small-scale analyses.
III.2 Spectroscopic redshift determination
@@
At each redshift step, we calculate the as in Equation 6 and find the redshift that minimizes this profile, which indicates that including an LAE component at that redshift improves the fit the most. The also indicates a level of confidence or SNR in our detection; higher indicates a more robust detection of an LAE. To achieve a more precise redshift estimate, we do an additional, finer scan along 5 pixels around this best redshift value in steps of 0.1 pixels, and assign the target the minimum redshift from this final scan. Averaged over the 881 Subaru targets, it took seconds to determine the redshift for one target spectrum.
@@ IV Results & Validation
IV.1 Redshift determination for Subaru LAE targets
We use the priors described in Section III to decompose and determine redshifts for DESI spectra of 881 LAE candidates targeted with Subaru that have VI redshifts between 2 and 4. We evaluate our method on a held-out test set, meaning that none of the spectra for which we show results were used to create the LAE prior covariance matrix.
Figure 4 shows the (calibrated as described in Section IV.2.3) as a function of redshift residual (new - VI ) for this test set. There is generally good agreement between the redshifts determined by our pipeline and those from visual inspection: 812 (92.3%) of the 881 evaluated spectra were determined by our method to have a redshift within 0.005 of the VI. Most disagreements between the two methods occur at low , which would indicate low SNR or an uncertain detection. Several factors could contribute to our pipeline determining redshifts outside of this threshold: some of the sources are Lyman Break Galaxies (LBGs) or other objects which do not have a strong Ly emission line; other objects that show strongly double-peaked Ly, or very weak lines; and some could be erroneous redshifts reported by the visual inspection.
To create a binary classifier (LAE or not LAE), one could impose a threshold value, above which any target is classified as being an LAE. Figure 6 shows the purity and completeness of the LAE sample as a function of the potential calibrated threshold value, where a “correct” detection is a target that has a pipeline redshift within 0.005 of its VI redshift. Note that we are evaluating our method on a pre-selected sample of solely LAE targets with no known non-LAEs, so a high purity across all thresholds is to be expected. Figures 4 and 6 suggest that a value of 25 could be an effective threshold value for classifying an object as an LAE, which roughly corresponds to a 5 cut.
In addition to the value, which denotes the improvement in the fit from adding an LAE component, the , hereafter referred to as the “total ”, indicates how good of a fit the entire model is to the data. Targets that are not LAEs but have other emission lines, such as additional sky residual lines or low-redshift interlopers, may have a large value, but will also likely have a large total value since our LAE component does not represent their emission line well. This total value can be used to identify potential false positives.
Figure 5 shows the and total values for the 881 Subaru LAE targets, colored by redshift residuals. For this dataset, there are five outliers with total . Visual inspection of these objects shows that they include quasar contamination and/or potential cosmic rays. Of these five outliers, our pipeline still identifies a redshift within 0.005 of the VI redshift for two of them.
As seen in Figure 4, the majority of poor redshift fits can be identified by imposing a cut. However, false positives can be further identified by imposing a total cut, drawing a vertical line on Figure 5. Additionally, a diagonal line can be drawn in this space to cleanly separate out the majority of redshift failures. A cut based on the total is likely useful for future applications of this method to larger, less pure datasets.
IV.2 Injection-recovery tests
IV.2.1 Uniform noise